OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 10x + 25 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

Please… factor this.

OpenStudy (oaktree):

Well, you could factor, OR you could use the discriminant, which is simpler.

OpenStudy (oaktree):

The discriminant is just the thing inside the square root in the quadratic formula, or b^2-4ac.

OpenStudy (anonymous):

how do i use the discriminant?

OpenStudy (oaktree):

If your discriminant (D) is 0, then you have 2 identical solutions.

OpenStudy (oaktree):

If it's negative, you have two different solutions that are IMAGINARY.

OpenStudy (oaktree):

If it's positive but a square number, then it's two different rational solutions.

OpenStudy (oaktree):

And lastly, if it's positive but not square-rootable, then it's two different irrational solutions.

OpenStudy (oaktree):

Really, it makes sense. If you think about it.

OpenStudy (oaktree):

Anyway, b^2-4ac = 10^2 - 4(1)(25) = 100-100=0.

OpenStudy (oaktree):

OOH! The Discriminant is 0! That means that you have...?

OpenStudy (anonymous):

2 identical solutions

OpenStudy (anonymous):

@OakTree is correct. The discriminant is faster. But you need to learn to factor, and that is an art (also a lot more fun).

OpenStudy (oaktree):

Good.

OpenStudy (oaktree):

The discriminant is a very useful tool, but remember that it won't tell you what the actual roots are. For that, you need the actual quadratic formula.

OpenStudy (anonymous):

Or factoring. :)

OpenStudy (anonymous):

okay thank you!